OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1

OpenStudy (oaktree):

If the slope is -1, you know the equation is of the form y = -x+b.

OpenStudy (oaktree):

So, now you only need to plug in the numbers from (1,-5).

OpenStudy (oaktree):

Plugging in, we have -5 = -(1) + b. This gives b=-6.

OpenStudy (oaktree):

So the equation is y=-x-6, so, in standard form, this is x+y=-6.

OpenStudy (oaktree):

Okay?

OpenStudy (oaktree):

?

OpenStudy (oaktree):

Anything unclear?

OpenStudy (anonymous):

i got it thanks